OpenStudy (anonymous):

4b^2-64 -------- -4b+16

OpenStudy (anonymous):

x =(-8-√-256)/2=-4-8i= -4.0000-8.0000i x =(-8+√-256)/2=-4+8i= -4.0000+8.0000i